OpenStudy (anonymous):

The point Seven comma three would not fall on the line for which of the following representations? Choices: A) 3x= -y+16 B) 2y-8=-x+5 C) 11-5x=-8y D) -4x+1=-9y I will give medals.

OpenStudy (whydoihavetosignup1):

ok i won't

OpenStudy (anonymous):

seven comma three= (7,3)

OpenStudy (anonymous):

plug in (x, y) = (7, 3) in each one and see if it holds true. for the first you want to see if 3x + y = 16, which is 3(7) + 3 = 16. this is false; 24 =/= 16/ A is the answer.

OpenStudy (anonymous):

So A is the correct answer?

OpenStudy (anonymous):

third one: does 11 = -8(3) + 5(7) = -24 + 35. sure does ^_^. therefore it falls on the line described by C. Yes. A is the answer.

OpenStudy (anonymous):

Oh I understand now thanks!!!